OpenStudy (anonymous):

im having a problem understanding algebra stuff. i can grasp how to multiply radicals. help anyone?

OpenStudy (amistre64):

\[\sqrt{a}~\sqrt{b}=\sqrt{a*b}\] not too sure how thats confusing ....

OpenStudy (anonymous):

i mean its hold on... https://media.glynlyon.com/g_alg02_2012/7/q11174.gif

OpenStudy (amistre64):

conjugates help with simplification

OpenStudy (anonymous):

rationalizing the denominator

OpenStudy (anonymous):

i get how to find the gonjugates.. maybe im just over thinking it -_-

OpenStudy (amistre64):

change the operation, and multiply

OpenStudy (amistre64):

(a+b) <-> (a-b) are conjugates

OpenStudy (amistre64):

the beauty is:\[(a+b)(a-b) = a^2 -b^2\]

OpenStudy (amistre64):

the top will just have to be whatever it plays out as ..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what i dont get is would it be 12 radical 36?

OpenStudy (amistre64):

\[\frac{c+d}{a-b}\frac{a+b}{a+b}=\frac{ca+cb+da+db}{a^2-b^2}\] the specific values are just distractions

OpenStudy (amistre64):

12(6) +9sqrt(12)+20sqrt(12)+15(4) ------------------------------ 16(6) - 9(4)

OpenStudy (amistre64):

uhh, got a little carried away :) *-9(2) is that under part :)

OpenStudy (amistre64):

and 15(2) up top ....

OpenStudy (amistre64):

12(6) +9sqrt(12)+20sqrt(12)+15(2) ------------------------------ 16(6) - 9(2) thats better

OpenStudy (anonymous):

*confused*

OpenStudy (amistre64):

not too sure what the confusion it ...

OpenStudy (anonymous):

wait.. hold on

OpenStudy (anonymous):

so 15 radical 2 is n top and -9 radical 2 on bottom. or is that supposed to be squared

OpenStudy (amistre64):

you recall how to "FOIL"? i dont really like that term but it seems to be what passes of math these days

OpenStudy (amistre64):

\((3\sqrt6+5\sqrt2)\) \((4\sqrt6+3\sqrt{2})\) ------------------------ \((4\sqrt6-3\sqrt{2})\) \((4\sqrt6+3\sqrt{2})\) \(3(4)\sqrt{6(6)}+5(4)\sqrt{2(6)}+3(3)\sqrt{6(2)}+5(3)\sqrt{2(2)}\) --------------------------------------------- \(4(4)\sqrt{6(6)}-3(4)\sqrt{2(6)}+3(4)\sqrt{2(6)}-3(3)\sqrt{2(2)}\) \(12\sqrt{36}+20\sqrt{12}+9\sqrt{12}+15\sqrt{4}\) ----------------------------- \(16\sqrt{36}-9\sqrt{4}\) \(12(6)+20\sqrt{12}+9\sqrt{12}+15(2)\) ----------------------------- \(16(6)-9(2)\)